Ops/Comm Scheduling meeting : 4/7/2017

Location : CR21 at 14:00 pm

Phone access : New phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

  • Actions
    • Hall Manager - water system change-over procedure (stands) - have prescription and draft document (Colin/Josef), stands
    • Progress on Hall list attachment:MICE-Work-List-May-2017.pdf (Colin) Last remaining item is believed to be verifying the interlocks; hope to get Trevor Hartnett down in next few weeks?
    • Status of miceraid5 disk cleanup (stands) EOverton now has a script but needs to run it: needs to hand over the procedure and script to someone, i.e. Durga Is Durga aware of this item?
    • Investigation of the widespread Watchdog Timer failure. CM look at substation monitor. Agit suggests writing a recovery document and ensuring we have most beneficial fail-safe and cold start conditions.
  • DC update
  • Hydrogen system update and plans
  • Cycle 2017 Issues : Status of prioritisation. Too long for this meeting - perhaps schedule a targetted discussion just on this issue?
  • Status of Support for Technical Endeavours And Knowledge (CWhyte)
  • AOB


  • Ed and Joe setting up new target computer in R78
  • Tracker group to source more scroll pumps. AB arrives 9th August
  • Air-conditioning in Hall is having issues: one of the units froze up completely.
    Paul Masterson is getting someone in.
  • One of the air-conditioning units in RR2 has failed. It's OK just on the back-up for now.
    A different company deals with this one
  • Ajit and Henry have worked through the issues list: C&M and Beamline issues have been identified and assigned.
    Steve Boyd will go through the Other issues with Colin Whyte next week (action).
  • Problem with DS. HN and MT investigating
  • Status of Support for Technical Endeavours And Knowledge (CWhyte)

Updated by Nebrensky, Henry almost 7 years ago · 6 revisions